# DART: A Diffusion-Based Autoregressive Motion Model for Real-Time Text-Driven Motion Control

### v1

@@ 3.2 DART: A Diffusion-Based Autoregressive Motion Primitive Model
Latent Motion Primitive Diffusion. Building on the motion representation provided by our trained motion primitive encoder $\mathcal{E}$, we transform the motion primitives from the raw feature space to a more refined latent space that is more suitable for generative learning with diffusion models. This latent motion primitive space is then used to train a diffusion denoiser for generating motion primitives conditioned on text prompts. The denoiser architecture is shown in Fig. 1. Given the noised latent variable $z_{t}$, the diffusion step $t$, the history motion $\mathbf{H}$, and a text prompt $c$, we train a transformer-based latent denoiser model $\mathcal{G}$ to predict the cleaned motion primitive latent variable $\hat{z}_{0}=\mathcal{G}(z_{t},t,\mathbf{H},c)$. The diffusion step is embedded using a small MLP, while the text prompt is encoded using the CLIP (Radford et al., 2021) text encoder. The text prompt is randomly masked out by a probability of 0.1 during training to enable classifier-free guidance (Ho & Salimans, 2022) during generation. During training, the motion primitive latent variable $z_{0}=\mathcal{E}(\mathbf{H},\mathbf{X})$ can be efficiently obtained with a single inference of the fixed encoder $\mathcal{E}$. The cleaned latent variable $\hat{z}_{0}$ can be converted back to the future frames using the fixed decoder $\mathcal{D}$: $\hat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\hat{z}_{0})$. The latent denoiser $\mathcal{G}$ is trained with the simple objective (Ho et al., 2020). Additionally, we use the reconstruction loss $L_{rec}$ and auxiliary losses $L_{aux}$ on the decoded future frames $\hat{\mathbf{X}}$, the same as training the motion primitive VAE. Notably, we use DDPM (Ho et al., 2020) with only 10 diffusion steps during both the training and inference of the denoiser model. This is sufficient for generating realistic samples due to the simplicity of the short motion primitives and enables highly efficient online generation. Moreover, we use the scheduled training (Ling et al., 2020; Bengio et al., 2015; Rempe et al., 2021) to progressively introduce the test-time distribution of the history motion $\mathbf{H}$, which improves the stability of long sequence online generation and the text prompt controllability for unseen poses. We refer to Appendix D for the details.
With the trained motion primitive decoder $\mathcal{D}$, latent denoiser $\mathcal{G}$ and a diffusion sampler $\mathcal{S}$ such as Ho et al. (2020) and Song et al. (2022), we can autoregressively generate motion sequences given the history motion seed $\mathbf{H}_{seed}$ and the online sequence of text prompts $C$, as shown in Alg. 1. During sampling, we use classifier-free guidance (Ho & Salimans, 2022) on the text condition with a guidance scale $w$:
$\mathcal{G}_{w}(z_{t},t,\mathbf{H},c)=\mathcal{G}(z_{t},t,\mathbf{H},\emptyset)+w\cdot(\mathcal{G}(z_{t},t,\mathbf{H},c)-\mathcal{G}(z_{t},t,\mathbf{H},\emptyset))$ (1)
Using the rollout algorithm, DART generates over 300 frames per second using a single RTX 4090 GPU, enabling real-time applications and online reinforcement-learning control as in Sec. 3.3.
@@
Optional Input: Latent noises $Z_{T}=[z_{T}^{1},...,z_{T}^{N}]$
@@
     sample noise $z_{T}^{i}$ from $\mathcal{N}(0,1)$ if not inputted
     $\hat{z}_{0}^{i}\leftarrow\mathcal{S}(\mathcal{G},z_{T}^{i},T,\mathbf{H},c^{i},w$) $\triangleright$ diffusion sample loop with classifier-free guidance
     $\hat{\mathbf{X}}\leftarrow\mathcal{D}(\mathbf{H},\hat{z}_{0}^{i})$
@@
3.3 Controllable Motion Synthesis via DART
Text-conditioned motion generation offers a user-friendly interface for controlling motions through natural language. However, relying solely on text limits precise spatial control, such as walking to a specific location or sitting in a designated spot. Therefore, it is necessary to incorporate motion control mechanisms to achieve precise spatial goals, including reaching a keyframe body pose, following joint trajectories, and interacting with scene objects.
We formulate the motion control task as generating the motion sequence $\mathbf{M}$ that minimizes its distance to a given spatial goal $g$ under a criterion function $\mathcal{F}(\cdot,\cdot)$ and the regularization from the scene and physical constraints $cons(\cdot)$:
$\mathbf{M}^{*}={\arg\!\min}_{\mathbf{M}}\mathcal{F}(\Pi(\mathbf{M}),g)+cons(\mathbf{M})$ (2)
where $g$ is the task-dependent spatial goal, e.g., $g$ is a keyframe body for the motion in-between task or the goal location for the navigation task. $\Pi(\cdot)$ is the goal-dependent projection function that maps motion sequence to the goal observation space, $cons(\cdot)$ denotes physical constraints such as avoiding collisions into scenes.
Directly solving the motion control task in the raw motion space often results in unrealistic motions since most samples in the raw motion space do not represent plausible motions. To improve the generated motion quality, many previous methods teacle such motion control tasks in a latent motion space, where samples can be mapped to plausible motions (Karunratanakul et al., 2024; Ling et al., 2020; Peng et al., 2022; Holden et al., 2015). DART offers a powerful text-conditioned latent motion space for such latent space control. DART learns a powerful generative model that can generate diverse and realistic motions from standard Gaussian samples $Z_{T}$. Using the deterministic DDIM (Song et al., 2022) sampler, we adapt DART sampling to function as a deterministic mapping from latent noises $Z_{T}$ to plausible motions. Using the latent noises $Z_{T}$ as the latent space, we transform the motion control task in Eq. 2 to latent space control as follows: Given the initial motion history $\mathbf{H}_{seed}$, a sequence of text prompts $C$, and a deterministic diffusion sampler $\mathcal{S}$, the rollout function in Alg. 1 can deterministically map a list of motion primitive latent noises $Z_{T}=[z_{T}^{1},..,z_{T}^{N}]$ to a motion sequence $\mathbf{M}$ by :
$\mathbf{M}=\textsc{rollout}(Z_{T},\mathcal{D},\mathcal{G},\mathbf{H}_{seed},C,T,w,\mathcal{S})$ (3)
We abbreviate it as $\textsc{rollout}(Z_{T},...)$ for short since $Z_{T}$ is the only variable in latent control. The minimization objective is converted as :
$Z_{T}^{*}={\arg\!\min}_{Z_{T}}\mathcal{F}(\Pi(\textsc{rollout}(Z_{T},...),g)+cons(\textsc{rollout}(Z_{T},...))$ (4)
@@ 3.3 Controllable Motion Synthesis via DART
Motion Control via Latent Diffusion Noise Optimization. One straightforward solution to this minimization problem (Eq. 4) is to directly optimize the latent noises $Z_{T}$ given the criterion function using gradient descent methods. The latent noise optimization is illustrated in Alg. 2.
@@
Input: Latent noises $Z_{T}=[z_{T}^{1},...,z_{T}^{N}]$, Optimizer $\mathcal{O}$, learning rate $\eta$, criterion $\mathcal{L}$, goal $g$ (for brevity, we do not reiterate the inputs of the rollout function defined in Alg. 1)
Output: motion sequence $\mathbf{M}$
@@
     $\mathbf{M}\leftarrow\textsc{rollout}(Z_{T},...)$
     $\nabla\leftarrow\nabla_{Z_{T}}\mathcal{L}(\mathbf{M})$
     $Z_{T}\leftarrow\mathcal{O}(Z_{T},\nabla/\|\nabla\|,\eta)$triangleright update using normalized gradient
@@
return $\mathbf{M}\leftarrow\textsc{rollout}(Z_{T})$
This optimization-based control framework is general and applicable for various spatial control tasks. We instantiate the latent noise optimization method in two example control scenarios: in-between motion generation and human-scene interaction generation. First, we addresss the motion in-between task that aims to generate the motion frames transition between given history and goal keyframes $g$ that is $f$ frames away conditioned on the text prompt $c$. We use the distance between the $f$-th frame of the generated motion and the goal keyframe as the optimization objective.
Second, we show that physical and scene constraints $cons(\cdot)$ can be incorporated to synthesize human motions in a contextual environment. Given the text prompts $C$ and spatial goal $g$ of the interaction anchor joint locations, e.g., locations of the pelvis when sitting, the objective is to generate motions that not only perform the desired interaction but also achieve the goal joint positions while adhering to scene constraints. We consider two scene constraints penalizing the human-scene collision and foot-floating artifacts, as detailed in Appendix. F.
@@
Figure 2: Architecture of the reinforcement learning-based control policy. The frozen DART diffusion denoiser and decoder models transform the latent actions into motion frames. The last predicted frames are canonicalized and provided to the policy model as the next step history condition.
Our latent motion primitive model naturally fits in the Markov decision process due to its autoregressive nature. We use the latent noises $z_{T}$ as the latent action space and train goal-conditioned policy models as controllers. The policy architecture is shown in Fig. 2. Our policy model employs the actor-critic (Sutton & Barto, 1998) architecture and is trained with the PPO (Schulman et al., 2017) algorithm. The state $\mathbf{s}^{i}$ includes the history motion observation $\mathbf{H}^{i}$, goal observation $g^{i}$, scene observation $s^{i}$, and the CLIP embedding of the text prompt $c^{i}$. The policy model takes in [$\mathbf{H}^{i}$, $g^{i}$, $s^{i}$, $c^{i}$] to predict the latent noise $z_{T}^{i}$ as the action. The latent noise $z_{T}^{i}$ is mapped to the future motion frames $\mathbf{X}^{i}$ using the frozen latent denoiser $\mathcal{G}$ and motion primitive decoder $\mathcal{D}$. The new history motion is extracted from the last $H$ predicted frames and fed to the policy network in the next step. We reformulate the minimization problem in Eq. 4 as reward maximization to train the policy.
We instantiate the reinforcement learning control with the text-conditioned goal-reaching task. Given a text prompt $c$ and a 2D goal location $g$, we aim to control the human to reach the goal location using the action specified by the text. The goal location is converted to a local observation consisting of its distance to the human pelvis, and its local direction in the human-centric coordinates frame. We consider a simple flat scene and the scene observation is the relative floor height to the human pelvis. The policy is trained with distance rewards encouraging the human pelvis to reach the goal location and scene constraint rewards penalizing foot skating and floor penetration. Further details can be found in appendix G. With the trained control policies, we can efficiently control a human to reach dynamic goals using specified skills like walking or hopping.
@@ 4 Experiments
We provide extensive experiments showing how DART can serve as a general model for text-conditioned temporal motion composition (4.1) and various motion generation tasks requiring precise spatial control via latent noise optimization (4.2) and reinforcement learning policy (4.3). The qualitative results and comparisons are attached in the supplementary video.
@@ 4.1 Text-Conditioned Temporal Motion Composition
We evaluate the generation results using metrics proposed in Guo et al. (2022) and Barquero et al. (2024). For each action segment, we evaluate the similarity between generation and dataset (FID), motion-text semantic alignment (R-prec, MM-DIST), and generation diversity (DIV). To evaluate smooth transitions between two segments, we measure the jerk (the derivative of acceleration) of the 30-frame transition clip centered at the splitting point of two action segments, reporting the peak jerk (PJ) and Area Under the Jerk (AUJ). Moreover, we profile all methods in a benchmark of generating one 5000-frame-long sequence and report the generation speed, the latency of getting the first generated frames, and memory usage. We also conduct binary user studies to evaluate the motion realism and motion-text semantic alignment of generation results, during which the participants are given generation results from two different methods and are asked to select the generation that is perceptually more realistic or better aligns with the action text stream in subtitles. We compare DART with baselines inlcuding TEACH (Athanasiou et al., 2022), DoubleTake (Shafir et al., 2024), a history-conditioned modification of T2M-GPT (Zhang et al., 2023a)(denoted as T2M-GPT*), and the state-of-the-art offline motion composition method FlowMDM (Barquero et al., 2024).
We present the quantitative results in Tab. 1 and Tab. 2. DART achieves the best FID in both the segment and transition evaluation, indicating the highest similarity to the dataset and best motion realism. DART also displays second-best jerk metrics indicating smooth action transitions. We observe that DART performs slightly worse than FlowMDM in motion-text semantic alignment (R-prec and MM-Dist) because of the online generation nature of DART. Natural action transitions require time to transit to the new action after receiving the new action prompt, leading to a delay in the emergence of the new action semantics. For instance, a human cannot immediately transition from kicking a leg in the air to stepping backward without first recovering to a standing pose. This inherent delay in transitions results in a motion embedding shift that impacts the R-prec metric of DART. In contrast, the offline baseline FlowMDM generates the entire sequence as a whole and requires oracle information of the full timeline of action segments to modulate compatibility between subsequent segments. The slight but natural action transition delay of DART is perceived as natural by humans, as shown in the user study results in Tab. 2 where DART is preferred over all the baselines, including FlowMDM, for both motion realism and motion-text semantic alignment in human evaluations.
@@
Table 2: Binary user study results comparing our method against baselines in generation realism and motion-text semantic alignment on text-conditioned temporal motion composition. We report the percentage of each method being voted better than the other (Ours vs. Baselines).

### v2

@@ 3.2 DART: A Diffusion-Based Autoregressive Motion Primitive Model
The denoiser model architecture is shown in Fig. 1. The diffusion step $t$ is embedded using a small MLP, while the text prompt $c$ is encoded using the CLIP (Radford et al., 2021) text encoder. The text prompt is randomly masked out by a probability of 0.1 during training to enable classifier-free guidance (Ho & Salimans, 2021) during generation. The cleaned latent variable $\hat{\mathbf{z}}_{0}$ can be converted back to the future frames using the frozen decoder $\mathcal{D}$: $\hat{\mathbf{X}}=\mathcal{D}(\mathbf{H},\hat{\mathbf{z}}_{0})$. We train the latent denoiser $\mathcal{G}$ using the simple objective (Ho et al., 2020). Additionally, we apply the reconstruction loss $L_{rec}$ and auxiliary losses $L_{aux}$ on the decoded future frames $\hat{\mathbf{X}}$. Notably, we use only 10 diffusion steps for both training and inference. This small number suffices for realistic sample generation due to the simplicity of our motion primitive representation, enabling highly efficient online generation. Moreover, we use the scheduled training (Ling et al., 2020; Bengio et al., 2015; Rempe et al., 2021) to progressively introduce the test-time distribution of the history motion $\mathbf{H}$, which improves the stability of long sequence online generation and the text prompt controllability for unseen poses. We refer to Appendix D for the details. With the trained motion primitive decoder $\mathcal{D}$, latent denoiser $\mathcal{G}$ and a diffusion sampler $\mathcal{S}$ such as DDPM and DDIM (Ho et al., 2020; Song et al., 2021a), we can autoregressively generate motion sequences given the history motion seed $\mathbf{H}_{seed}$ and the online sequence of text prompts $C$, as shown in Alg. 1. During sampling, we use classifier-free guidance (Ho & Salimans, 2021) on the text condition with a guidance scale $w$: $\mathcal{G}_{w}(\mathbf{z}_{t},t,\mathbf{H},c)=\mathcal{G}(\mathbf{z}_{t},t,\mathbf{H},\emptyset)+w\cdot(\mathcal{G}(\mathbf{z}_{t},t,\mathbf{H},c)-\mathcal{G}(\mathbf{z}_{t},t,\mathbf{H},\emptyset)).$ Using the rollout algorithm, DART generates over 300 frames per second using a single RTX 4090 GPU, enabling real-time applications and online reinforcement-learning control as in Sec. 3.3.
@@
Optional Input: Latent noises $\mathbf{Z}_{T}=[\mathbf{z}_{T}^{1},...,\mathbf{z}_{T}^{N}]$
@@
     sample noise $\mathbf{z}_{T}^{i}$ from $\mathcal{N}(0,1)$ if not inputted
     $\hat{\mathbf{z}}_{0}^{i}\leftarrow\mathcal{S}(\mathcal{G},\mathbf{z}_{T}^{i},T,\mathbf{H},c^{i},w$) $\triangleright$ diffusion sample loop with classifier-free guidance
     $\hat{\mathbf{X}}\leftarrow\mathcal{D}(\mathbf{H},\hat{\mathbf{z}}_{0}^{i})$
@@
3.3 Spatially Controllable Motion Synthesis via DART
Text-conditioned motion generation offers a user-friendly interface for controlling motions through natural language. However, relying solely on text limits precise spatial control, such as walking to a specific location or sitting in a designated spot. Therefore, it is necessary to incorporate motion control mechanisms to achieve precise spatial goals, including reaching a keyframe body pose, following joint trajectories, and interacting with scene objects. We formulate the motion control task as generating the motion sequence $\mathbf{M}$ that minimizes its distance to a given spatial goal $g$ under a criterion function $\mathcal{F}(\cdot,\cdot)$ and the regularization from the scene and physical constraints $cons(\cdot)$:
$\mathbf{M}^{*}={\arg\!\min}_{\mathbf{M}}\mathcal{F}(\Pi(\mathbf{M}),g)+cons(\mathbf{M}),$ (2)
where $g$ is the task-dependent spatial goal (e.g., a keyframe body for motion in-between tasks or a target location for navigation tasks), $\Pi(\cdot)$ is the projection function that extracts goal-relevant features from motion sequences and maps them into the task-aligned observation space, and $cons(\cdot)$ denotes physical and scene constraints, such as preventing scene collisions and floating bodies.
Directly solving the motion control task in the raw motion space often results in unrealistic motions since most samples in the raw motion space do not represent plausible motions. To improve the generated motion quality, many previous methods tackle such motion control tasks in a latent motion space, where samples can be mapped to plausible motions (Karunratanakul et al., 2024b; Ling et al., 2020; Peng et al., 2022; Holden et al., 2015). DART offers a powerful text-conditioned latent motion space for such latent space control, as it learns a generative model capable of producing diverse and realistic motions from standard Gaussian samples. Using the deterministic DDIM (Song et al., 2021a) sampler, we adapt DART sampling to function as a deterministic mapping from latent noises $\mathbf{Z}_{T}$ to plausible motions. This allows us to reformulate the motion control task in Eq. 2 as a latent space control problem as follows: Given the initial motion history $\mathbf{H}_{seed}$, a sequence of text prompts $C$, the pretrained DART models, and a deterministic diffusion sampler $\mathcal{S}$, the rollout function in Alg. 1 can deterministically map a list of motion primitive latent noises $\mathbf{Z}_{T}=[\mathbf{z}_{T}^{1},..,\mathbf{z}_{T}^{N}]$ to a motion sequence conditioned on the history motion seed and text prompts: $\mathbf{M}=\textsc{rollout}(\mathbf{Z}_{T},\mathbf{H}_{seed},C).$ The minimization objective is converted as:
${\mathbf{Z}_{T}}^{*}={\arg\!\min}_{\mathbf{Z}_{T}}\mathcal{F}(\Pi(\textsc{rollout}(\mathbf{Z}_{T},\mathbf{H}_{seed},C)),g)+cons(\textsc{rollout}(\mathbf{Z}_{T},\mathbf{H}_{seed},C))$ (3)
@@ 3.3 Spatially Controllable Motion Synthesis via DART
Motion Control via Latent Diffusion Noise Optimization. One straightforward solution to this minimization problem (Eq. 3) is to directly optimize the latent noises $\mathbf{Z}_{T}$ given the criterion function using gradient descent methods (Kingma & Ba, 2015; Karunratanakul et al., 2024a). The latent noise optimization is illustrated in Alg. 2. This optimization-based control framework is general and applicable for various spatial control tasks. We instantiate the latent noise optimization method in two example control scenarios: in-between motion generation and human-scene interaction generation.
@@
Input: Latent noises $\mathbf{Z}_{T}=[\mathbf{z}_{T}^{1},...,\mathbf{z}_{T}^{N}]$, Optimizer $\mathcal{O}$, learning rate $\eta$, and goal $g$. (For brevity, we do not reiterate the inputs of the rollout function defined in Alg. 1 and criterion terms in Eq. 2)
Output: a motion sequence $\mathbf{M}.$
@@
     $\mathbf{M}\leftarrow\textsc{rollout}(\mathbf{Z}_{T},\mathbf{H}_{seed},C)$
     $\nabla\leftarrow\nabla_{\mathbf{Z}_{T}}(\mathcal{F}(\Pi(\mathbf{M}),g)+cons(\mathbf{M}))$
     $\mathbf{Z}_{T}\leftarrow\mathcal{O}(\mathbf{Z}_{T},\nabla/\|\nabla\|,\eta)$triangleright update using normalized gradient
@@
return $\mathbf{M}\leftarrow\textsc{rollout}(\mathbf{Z}_{T},\mathbf{H}_{seed},C)$
First, we address the motion in-between task that aims to generate the motion frames transition between given history and goal keyframes $g$ that is $f$ frames away conditioned on the text prompt $c$. We use the distance between the $f$-th frame of the generated motion and the goal keyframe as the optimization objective. Second, we show that physical and scene constraints $cons(\cdot)$ can be incorporated to synthesize human motions in a contextual environment. Given input 3D scenes, text prompts $C$, and spatial goals $g$ of the interaction anchor joint locations, e.g., locations of the pelvis when sitting, the objective is to generate motions that not only perform the desired interaction but also achieve the goal joint positions while adhering to scene constraints. During the optimization, the 3D scenes are represented as signed distance fields (SDF) to compute body-scene distances, which serve as the basis for deriving human-scene contact and collision metrics that encourage foot-floor contact and scene collision avoidance, as detailed in Appendix. F.
@@
Figure 2: Architecture of the reinforcement learning-based control policy. The pretrained DART diffusion denoiser and decoder models transform the latent actions into motion frames. The last predicted frames are canonicalized and provided to the policy model as the next step history condition.
Our latent motion primitive model naturally fits in the Markov decision process due to its autoregressive nature. We use the latent noises $\mathbf{z}_{T}$ as the latent action space and train goal-conditioned policy models as controllers. The policy architecture is shown in Fig. 2. Our policy model employs the actor-critic (Sutton & Barto, 1998) architecture and is trained with the PPO (Schulman et al., 2017) algorithm. The state $\mathbf{s}^{i}$ includes the history motion observation $\mathbf{H}^{i}$, goal observation $g^{i}$, scene observation $s^{i}$, and the CLIP embedding of the text prompt $c^{i}$. The policy model takes in [$\mathbf{H}^{i}$, $g^{i}$, $s^{i}$, $c^{i}$] to predict the latent noise $\mathbf{z}_{T}^{i}$ as the action. The latent noise $\mathbf{z}_{T}^{i}$ is mapped to the future motion frames $\mathbf{X}^{i}$ using the frozen latent denoiser $\mathcal{G}$ and motion primitive decoder $\mathcal{D}$. The new history motion is extracted from the last $H$ predicted frames and fed to the policy network in the next step. We reformulate the minimization problem in Eq. 3 as reward maximization to train the policy.
We instantiate the reinforcement learning control with the text-conditioned goal-reaching task. Given a text prompt $c$ and a 2D goal location $g$, we aim to control the human to reach the goal location using the action specified by the text. The goal location is transformed into a local observation, which includes its distance to the body pelvis at the last history frame and its local direction within the human-centric coordinate frame. We consider a simple flat scene and the scene observation is the relative floor height to the body pelvis at the first history frame. The policy is trained with distance rewards encouraging the human pelvis to reach the goal location and scene constraint rewards penalizing foot skating and floor penetration. Further details can be found in appendix G. With the trained control policies, we can efficiently control a human to reach dynamic goals using specified skills like walking or hopping.
@@ 4 Experiments
We provide extensive experiments showing how DART can serve as a general model for text-conditioned temporal motion composition (4.1) and various motion generation tasks requiring precise spatial control via latent noise optimization (4.2) and reinforcement learning policy (4.3). Qualitative results and comparisons are available in the supplementary videos on the project page.
@@ 4.1 Text-Conditioned Temporal Motion Composition
We evaluate the generation results using metrics proposed in Guo et al. (2022) and Barquero et al. (2024). For each action segment, we evaluate the similarity between generation and dataset (FID), motion-text semantic alignment (R-prec, MM-DIST), and generation diversity (DIV). To evaluate smooth transitions between two segments, we measure the jerk (the derivative of acceleration) of the 30-frame transition clip centered at the splitting point of two action segments, reporting the peak jerk (PJ) and Area Under the Jerk (AUJ). Moreover, we profile all methods in a benchmark of generating one 5000-frame-long sequence and report the generation speed, the latency of getting the first generated frames, and memory usage. We also conduct human preference studies to evaluate motion realism and motion-text semantic alignment of generation results, during which the participants are given generation results from two different methods and are asked to select the generation that is perceptually more realistic or better aligns with the action text stream in subtitles. We compare DART with baselines inlcuding TEACH (Athanasiou et al., 2022), DoubleTake (Shafir et al., 2024), a history-conditioned modification of T2M-GPT (Zhang et al., 2023a)(denoted as T2M-GPT*), and the state-of-the-art offline motion composition method FlowMDM (Barquero et al., 2024).
We present the quantitative results in Tab. 1 and Tab. 2. DART achieves the best FID in both the segment and transition evaluation, indicating the highest similarity to the dataset and best motion realism. DART also displays second-best jerk metrics indicating smooth action transitions. We observe that DART performs slightly worse than FlowMDM in motion-text semantic alignment (R-prec and MM-Dist) because of the online generation nature of DART. Natural action transitions require time to transit to the new action after receiving the new action prompt, leading to a delay in the emergence of the new action semantics. For instance, a human cannot immediately transition from kicking a leg in the air to stepping backward without first recovering to a standing pose. This inherent delay in transitions results in a motion embedding shift that impacts the R-prec metric of DART. In contrast, the offline baseline FlowMDM generates the entire sequence as a whole and requires oracle information of the full timeline of action segments to modulate compatibility between subsequent segments. The slight but natural action transition delay of DART is perceived as natural by humans, as shown in the preference study results in Tab. 2. DART is preferred over all the baselines, including FlowMDM, for both motion realism and motion-text semantic alignment in human evaluations.
@@
Table 2: Human preference study results comparing our method against baselines in generation realism and motion-text semantic alignment on text-conditioned temporal motion composition. We report the percentage of each method being voted better than the other (Ours vs. Baselines).